# AD-LLM: Benchmarking Large Language Models for Anomaly Detection

### v1

@@ 1 Introduction
Anomaly detection (AD) is an important topic in machine learning (ML) that identifies samples differing from the general distribution Zhao et al. (2019); Liu et al. (2024b). This ability is critical for many practical applications, such as fraud detection Abdallah et al. (2016), medical diagnosis Fernando et al. (2021), software engineering Sun et al. (2022), and industrial system monitoring Sun et al. (2023). Within natural language processing (NLP), AD is also important for finding unusual text instances, which is needed for detecting spam Rao et al. (2021), misinformation Islam et al. (2020), or unusual user behavior Xue et al. (2023).
@@
Key Takeaways. Our results reveal several noteworthy insights: (i) LLMs can achieve superior zero-shot AD performance, often outpacing conventional methods without relying on task-specific data; (ii) Enriching LLM inputs with additional context—such as anomaly category names or descriptive prompts—further boosts detection quality; (iii) Employing LLM-driven data augmentation enhances AD performance, though the effectiveness varies with model complexity and dataset properties; (iv) LLM-based model selection can approach top-performing baselines, but improving interpretability and providing dataset-specific rationales remains an open area. These suggest future work that systematically integrates external knowledge, refines prompt engineering, and develops strategies to ensure more transparent, context-sensitive LLM recommendations in AD tasks.
@@ 2.1 Related Work
Recent studies have explored the role of LLMs in AD, highlighting both opportunities and challenges. Xu and Ding (2024) proposes a taxonomy categorizing LLMs as either detection or generative tools, but their work lacks experimental benchmarks. Similarly, Jiang et al. (2024b) presents MMAD, a benchmark designed for industrial AD, focusing on image datasets yet limiting its applicability to other modalities. Liu et al. (2024a) evaluates LLMs like Llama for out-of-distribution (OOD) detection, demonstrating the effectiveness of cosine distance detectors with isotropic embeddings. However, their work primarily focuses on traditional metrics and lacks insights into advanced LLM capabilities such as data augmentation and zero-shot detection.
@@ 2.3 Common Experimental Settings
LLMs and Hardware. We select two LLMs as main backbones: (1) Llama 3.1 8B Instruct Dubey et al. (2024) and (2) GPT-4o OpenAI (2024a). For brevity, we refer to the “Llama 3.1 8B Instruct” as “Llama 3.1”. Llama 3.1 runs on NVIDIA RTX 6000 Ada, 48 GB RAM workstations. GPT-4o is accessed through Azure OpenAI API with the “2024-08-01-preview” version and OpenAI official API. Seed is set $=42$ for reproducibility. Specific experimental settings will be highlighted separately in each subsequent task.
Table 1: Performance comparison of LLM-based detectors and baseline methods across five datasets, evaluated under two settings as described in §3.2 with AUROC and AUPRC as the metrics (higher ($\uparrow$), the better). We list the best and the second-best baseline methods in each dataset. Complete results are provided in Appx. A2. The best results are highlighted in bold, and the second-best results are underlined. GPT-4o achieves the best performance consistently across all datasets, while Llama 3.1 also shows competitive performance, highlighting the effectiveness of LLM-based zero-shot anomaly detection.
@@ 3.2 Problem Statement and Designs
Given a test set $\mathcal{D}_{\text{test}}=\{x_{1},x_{2},\dots,x_{n}\}$ of text samples, where each sample $x_{i}$ belongs to either a normal category or an anomaly category, the objective is to identify the anomalous samples using a pre-trained LLM $f_{\text{LLM}}$ in a zero-shot setting without any task-specific training data.
Evaluation Protocol. We focus on the ability of LLMs to detect anomalies without additional training data. We consider two settings, each reflecting different levels of prior knowledge:
@@
$\mathcal{P}$ $=T(x_{i},\mathcal{C}_{\text{normal}},\mathcal{C}_{\text{anomaly}}^{*})$ (1)
$(r,s)$ $=f_{\text{LLM}}(\mathcal{P})$
Here, $T(\cdot)$ constructs the prompt $\mathcal{P}$ for a test sample $x_{i}$, including known category information. The anomaly category $\mathcal{C}_{\text{anomaly}}$ is included only in the “Normal + Anomaly” setting, denoted as $\mathcal{C}_{\text{anomaly}}^{*}$. The LLM $f_{\text{LLM}}$ processes the prompt to produce an anomaly score $s$ and an explanation $r$ that describes the reasoning. This setup allows a systematic evaluation of LLMs in zero-shot AD, using prompt-based inference to handle different levels of prior knowledge. See details in Appx. B.
@@ 3.3 Results, Insights, and Future Directions
We select Llama 3.1 and GPT-4o as zero-shot detectors with temperature $=0$ for stable outputs.
LLMs are effective in zero-shot AD, surpassing existing training-based AD algorithms. We compare Llama 3.1-based and GPT-4o-based zero-shot detectors with baseline methods across five datasets in Table 1. GPT-4o consistently outperforms all baselines, and Llama 3.1 often ranks near the top. Despite operating with limited prior information, LLMs exhibit significant potential for anomaly detection tasks. These results highlight the strength of LLMs in zero-shot AD scenarios.
Additional context helps. Table 1 shows improved AUROC and AUPRC when using both normal and anomaly categories (“Normal + Anomaly”) compared to using only normal categories (“Normal Only”). By providing more contextual information, LLMs better distinguish anomalous samples, increasing detection effectiveness. These results suggest that augmenting LLMs with richer information enhances their ability to detect anomalies.
Figure 2: LLM-based AD precision-recall comparison between two settings across five datasets. “Normal Only” and “Normal + Anomaly” are two settings discussed in §3.2. The anomaly score threshold $=0.5$ is used to decide the anomaly. Setting “Normal + Anomaly” outperforms setting “Normal Only” in most cases. In the “AG News” and “BBC News” datasets, there is a trade-off between precision and recall across two settings.
Trade-offs in detection focus. Figure 2 presents precision-recall comparisons for the two settings. Including anomaly information generally improves recall and sometimes precision. However, for “AG News” and “BBC News,” we observe a recall gain at the cost of precision. This indicates that adding anomaly details may shift the LLM’s detection balance, emphasizing one metric over another. Thus, one must carefully select the information provided to LLMs, as it significantly impacts their detection priorities and the balance among metrics.
Future Direction 1: Improve Context Integration. Providing additional context improves detection, as seen in “Normal + Anomaly.” Future work may involve more systematic ways to integrate domain-specific details based on the detection priority (e.g., precision vs. recall), such as prompt design or retrieval-augmented methods Gao et al. (2023).
Future Direction 2: Optimize for Real-world Deployment. Despite their effectiveness, LLM-based zero-shot AD is inherently time-consuming and costly Sinha et al. (2024). Reducing computational overhead is important for deploying LLMs in real settings, especially for AD applications, which are often time-critical. Methods like quantization Dettmers et al. (2023); Xiao et al. (2023), pruning Sun et al. (2024); Fu et al. (2024), and knowledge distillation Wang et al. (2024b); Fu et al. (2023) can help reduce the model size and inference time while maintaining good performance.
@@ 4.2 Generating Synthetic Samples for Training-based AD Models
Given a small training set $\mathcal{D}_{\text{small\_train}}=\{x_{1},x_{2},\dots,x_{m}\}$ of normal samples, the goal is to produce a synthetic dataset $\mathcal{D}_{\text{synth}}=\{\tilde{x}_{1},\tilde{x}_{2},\dots,\tilde{x}_{n}\}$ using a pre-trained LLM $f_{\text{LLM}}$. The combined dataset $\mathcal{D}_{\text{DA}}=\mathcal{D}_{\text{small\_train}}\cup\mathcal{D}_{\text{synth}}$ is used to train an unsupervised AD method $M$, improving performance compared to using $\mathcal{D}_{\text{small\_train}}$ alone.
Evaluation Protocol. To evaluate the impact of LLM-generated synthetic data, we set unsupervised AD baselines listed in §A.2 in a scenario with limited training data. LLMs are then utilized to generate a synthetic training dataset. However, direct prompting often leads to highly repetitive outputs, even with high decoding temperatures Long et al. (2024). Additionally, LLMs face constraints such as token limits (e.g., GPT-4’s maximum output of 4,096 tokens) and challenges in processing long contexts Gao et al. (2024). To address these issues, we adopt a multi-step strategy:
  • Step1: Keyword Generation: Generate groups of keywords in one inquiry. Each group consists of three keywords w/ different levels of granularity: broad/general, intermediate, and fine-grained.
@@
By separating keyword generation from sample creation and enforcing different granularity levels, this approach introduces controlled variability and thematic breadth without causing the model to produce overly lengthy or repetitive data. The resulting synthetic samples are more likely to be contextually rich and semantically diverse. Further details are provided in Appx. C.1
Figure 3: Performance comparison of AD baselines trained with and without LLM-generated synthetic data across five datasets. The use of LLM-generated synthetic data consistently enhances AD performance, with improvements varying by algorithm.
Table 2: Performance (and $\triangle$ changes) of LLM-based detectors with augmented descriptions under two settings in §3.2. The description generators and LLM-based detectors adopt the same LLM backbone. Values in brackets indicate changes compared to the results in Table 1. green denotes for improvements and red for declines. Changes below $0.03$ are not colored, reflecting minor fluctuations. Augmented descriptions improve performance in most cases, particularly in AUROC, showcasing their effectiveness in enhancing LLM-based AD.
@@
Results, Insights, and Future Directions. We apply the augmentation strategy to improve two unsupervised AD methods identified as strong baselines from Task 1 results: “OpenAI + LUNAR” Goodge et al. (2022) and “OpenAI + LOF” Breunig et al. (2000). We use GPT-4o with a temperature of $1.0$ to introduce sufficient variability in synthetic samples.
LLM-generated synthetic data consistently improves AD performance. As shown in Figure 3, LLM-generated synthetic data provides overall performance improvements across anomaly detection methods. Notably, “OpenAI + LUNAR” demonstrates consistent gains across all five datasets, emphasizing the robustness of this approach. Significant improvements are observed in datasets like “SMS Spam” and “N24 News,” where limited original training data posed significant challenges for traditional approaches. These results suggest that LLM-generated data is particularly effective in addressing data scarcity, complementing models like “LUNAR” that benefit from robust feature representations and semantic reasoning.
Improvement depends on model complexity and data characteristics. While “OpenAI + LUNAR” sees steady improvements, “OpenAI + LOF”—a simpler density-based model—exhibits variable outcomes. In data-scarce scenarios such as “SMS Spam” and “N24 News,” synthetic data assists LOF. However, for datasets like “AG News” and “BBC News,” performance declines, possibly due to the synthetic samples shifting the underlying data distribution in ways that do not align with LOF’s density assumptions. These results suggest that the effectiveness of LLM-based augmentation can depend on both the algorithm’s complexity and the dataset’s intrinsic structure.
@@ 4.3 Generating Category Descriptions for LLM-based Detectors
Results, Insights, and Future Directions. We utilize Llama 3.1 and GPT-4o to generate category descriptions. To balance the diversity and precision of the generation, we set the temperature to $0.5$.
Augmented descriptions improve LLM-based AD. As shown in Table 2, incorporating category descriptions increases AUROC scores in most datasets. This suggests that the added semantic information helps LLM-based detectors discriminate anomalous samples more effectively, especially when the dataset’s inherent structure aligns well with the contextual signals embedded in the descriptions. For example, in datasets like “IMDB Reviews” and “BBC News,” providing richer textual representations of normal and anomalous classes translates to noticeable gains in both metrics.
Trade-offs arise between different metrics. While AUROC improvements are widespread, some datasets exhibit declines in AUPRC, indicating that enhanced semantic cues may not uniformly boost all aspects of detection. This trade-off could occur if broad, high-level descriptions cause certain anomalies to appear more similar to normal samples, reducing precision. For instance, in “N24 News” and “SMS Spam,” the provided descriptions might introduce overlapping attributes between categories, making it more challenging to maintain high precision. These results underscore the importance of calibrating the level and type of detail included in category descriptions to suit the specific dataset characteristics.
Future Direction 4: Select Representative Samples. An effective way to refine category descriptions is to ground them in representative samples from the dataset. Sampling strategies based on clustering Axiotis et al. (2024) or diversity maximization Moumoulidou et al. (2020) can identify prototype examples that guide LLMs toward producing more tailored and context-aware descriptions. By referencing these representative samples, future approaches may generate descriptions that better reflect subtle category distinctions, ultimately improving both ranking quality and the balance between precision and recall.
@@ 5.1 Motivation
LLMs offer a promising zero-shot alternative by utilizing their extensive pre-trained knowledge to analyze datasets and recommend suitable models without relying on past performance metrics Qin et al. (2024). They can streamline the model selection process, reducing manual overhead and domain knowledge requirements while also improving adaptability to novel data scenarios.

### v3

@@ 1 Introduction
Anomaly detection (AD) is an important topic in machine learning (ML) that identifies samples differing from the general distribution Zhao et al. (2019); Liu et al. (2024c). This ability is critical for many practical applications, such as fraud detection Abdallah et al. (2016), medical diagnosis Fernando et al. (2021), software engineering Sun et al. (2022), and industrial system monitoring Sun et al. (2023). Within natural language processing (NLP), AD is also important for finding unusual text instances, which is needed for detecting spam Rao et al. (2021), misinformation Islam et al. (2020), or unusual user behavior Xue et al. (2023).
@@
Key Takeaways. Our results reveal several noteworthy insights: (i) LLMs can achieve superior zero-shot AD performance, often outperforming conventional methods without relying on task-specific data. (ii) Enriching LLM inputs with additional context, such as anomaly category names or descriptive prompts, further boosts detection quality. (iii) Employing LLM-driven data augmentation enhances AD performance, though the effectiveness varies with model features and dataset properties. (iv) LLM-based model selection can approach top-performing baselines, but improving interpretability and providing dataset-specific rationales remains an open area. These suggest future work that systematically integrates external knowledge, refines prompt engineering, and develops strategies to ensure more transparent, context-aware LLM recommendations in AD tasks.
@@ 2.1 Related Work
Recent studies have explored the role of LLMs in AD, highlighting both opportunities and challenges. Xu and Ding (2024) proposes a taxonomy categorizing LLMs as either detection or generative tools, but their work lacks experimental benchmarks. Similarly, Jiang et al. (2024b) presents MMAD, a benchmark designed for industrial AD, focusing on image datasets yet limiting its applicability to other modalities. Liu et al. (2024b) evaluates LLMs like Llama for out-of-distribution (OOD) detection, demonstrating the effectiveness of cosine distance detectors with isotropic embeddings achieved from LLMs. However, their study does not explore advanced LLM capabilities like data augmentation and zero-shot detection.
@@ 2.3 Common Experimental Settings
LLMs and Hardware. We select three LLMs as main backbones: (1) Llama 3.1 8B Instruct (referred to as Llama 3.1) (Dubey et al., 2024), (2) GPT-4o (OpenAI, 2024a), and (3) DeepSeek-V3 (Liu et al., 2024a). Llama 3.1 represents an open-source model with accessible size and cost, GPT-4o serves as a closed-source model with advanced capabilities, and DeepSeek-v3 represents the Mixture-of-Experts (MoE) architecture.
Llama 3.1 runs on one NVIDIA RTX 6000 Ada GPU with 48 GB RAM. GPT-4o and DeepSeek-V3 are accessed through official APIs. Seed is set $=42$ for reproducibility. Specific experimental settings are highlighted separately in each subsequent task.
Table 1: Performance comparison of LLM-based detectors and baseline methods across five datasets, evaluated under two settings as described in §3.2 with AUROC and AUPRC as the metrics (higher ($\uparrow$), the better). Complete results are provided in Appx. A2. The best results are highlighted in bold, and the second-best results are underlined.
@@
DeepSeek-V3
(1) with $\mathcal{C}_{\text{normal}}$ 0.9104 0.6442 0.8206 0.5604 0.8544 0.6808 0.8207 0.4495 0.8797 0.5963
(2) with $\mathcal{C}_{\text{normal}}$, $\mathcal{C}_{\text{anomaly}}$ 0.9273 0.7817 0.9581 0.8972 0.9626 0.8569 0.9514 0.7730 0.9535 0.7914
@@ 3.2 Problem Statement and Designs
Given a test set $\mathcal{D}_{\text{test}}=\left\{x_{1},x_{2},\dots,x_{n}\right\}$ of text samples, where each sample $x_{i}$ belongs to either a normal category or an anomaly category, the objective is to identify the anomalous samples using a pre-trained LLM $f_{\text{LLM}}$ in a zero-shot setting without any task-specific training data.
Evaluation Protocol. We consider two settings, each reflecting different levels of prior knowledge:
@@
$\mathcal{P}$ $=T\left(x_{i},\mathcal{C}_{\text{normal}},\mathcal{C}_{\text{anomaly}}^{*}\right)$ (1)
$\left(r,s\right)$ $=f_{\text{LLM}}\left(\mathcal{P}\right)$
Here, $T(\cdot)$ constructs the prompt $\mathcal{P}$ for a test sample $x_{i}$, including known category information. The anomaly category is included only in the “Normal + Anomaly” setting, denoted as $\mathcal{C}_{\text{anomaly}}^{*}$. The LLM $f_{\text{LLM}}$ processes the prompt to produce a verbal anomaly score $s$ and an explanation $r$ that describes the reasoning. This setup allows a systematic evaluation of LLMs in zero-shot AD, using prompt-based inference to handle different levels of prior knowledge. See details in Appx. B.
@@ 3.3 Results, Insights, and Future Directions
We select Llama 3.1, GPT-4o, and DeepSeek-V3 as zero-shot detectors. Temperature is set as $=0$ for stable outputs.
LLMs are effective in zero-shot AD, surpassing existing training-based AD algorithms. We compare LLM-based zero-shot detectors with top baselines across five datasets in Table 1. GPT-4o and DeepSeek-V3 consistently outperform baselines; Llama 3.1 shows competitive performance when anomaly information is available. Despite operating with limited prior information, LLMs exhibit significant potential for anomaly detection tasks. These results highlight the strength of LLMs in zero-shot AD scenarios.
Additional context helps. Table 1 shows that LLM-based detectors achieve improved AUROC and AUPRC when transitioning from setting “Normal Only”, which uses only $\mathcal{C}_{\text{normal}}$, to setting “Normal + Anomaly”, which includes both $\mathcal{C}_{\text{normal}}$ and $\mathcal{C}_{\text{anomaly}}$. These results indicate that richer contextual information improves the LLMs’ ability to distinguish anomalous samples and enhances detection performance.
Future Direction 1: Improve Context Integration. Providing additional context improves detection, as seen in “with $\mathcal{C}_{\text{normal}}$, $\mathcal{C}_{\text{anomaly}}$.” Future work may involve more systematic ways to integrate domain-specific details, such as prompt design or retrieval-augmented methods Gao et al. (2023).
Future Direction 2: Optimize for Real-world Deployment. Despite their effectiveness, LLM-based zero-shot AD is inherently time-consuming and costly during the inference Sinha et al. (2024). Reducing computational overhead is important for deploying LLMs in real settings, especially for AD applications, which are often time-critical. Methods like quantization Dettmers et al. (2023); Xiao et al. (2023), pruning Sun et al. (2024); Fu et al. (2024), and knowledge distillation Wang et al. (2024b); Fu et al. (2023) can help reduce the model size and inference time while maintaining good performance.
@@ 4.2 Generating Synthetic Samples for Training-based AD Models
Given a small training set $\mathcal{D}_{\text{small\_train}}=\left\{x_{1},x_{2},\dots,x_{m}\right\}$ of normal samples, the goal is to produce a synthetic dataset $\mathcal{D}_{\text{synth}}=\left\{\tilde{x}_{1},\tilde{x}_{2},\dots,\tilde{x}_{n}\right\}$ using a pre-trained LLM $f_{\text{LLM}}$. The combined dataset $\mathcal{D}_{\text{DA}}=\mathcal{D}_{\text{small\_train}}\cup\mathcal{D}_{\text{synth}}$ is used to train an unsupervised AD method $M$, improving performance compared to using $\mathcal{D}_{\text{small\_train}}$ alone.
Evaluation Protocol. To evaluate the impact of LLM-generated synthetic data, we set unsupervised AD baselines listed in Appx. A.2 in a scenario with limited training data. LLMs are then utilized to generate a synthetic training dataset. However, direct prompting often leads to highly repetitive outputs, even with high decoding temperatures Long et al. (2024). Additionally, LLMs face constraints such as token limits and challenges in processing long contexts Gao et al. (2024). To address these issues, we adopt a multi-step strategy:
  • Step1: Keyword Generation: Generate groups of keywords in one inquiry. Each group contains three keywords with a different level of granularity: broad/general, intermediate, or fine-grained.
@@
Separating keyword generation from sample creation and enforcing different granularity levels ensures controlled variability and prevents overly long or repetitive outputs. This results in more contextually rich and diverse synthetic samples.
To scale up further, we generate synthetic data in multiple rounds. In each round, we adjust the random seed, decoding temperature, and prompt template to ensure diversity. Further details are provided in Appx. C.1
Figure 2: Average performance over five datasets of AD baselines trained on limited data, w/ or w/o LLM-generated synthetic data, and on full datasets across five datasets. (a) Detectors that benefit from augmentation. (b) Detectors that degrade with augmentation.
Table 2: Performance (and $\triangle$ changes) of LLM-based detectors with augmented descriptions under two settings in §3.2. The description generators and LLM-based detectors adopt the same backbone. Values in brackets indicate changes compared to the results in Table 1. Green denotes for improvements and red for declines. Changes below $0.03$ are not colored for better visualization, also reflecting minor fluctuations.
@@
DeepSeek-V3
(1) with $\mathcal{C}_{\text{normal}}$ 0.8791 0.5180 0.8800 0.6170 0.9612 0.7888 0.8261 0.3949 0.9262 0.6128
(-0.0482) (-0.1262) (+0.0594) (+0.0566) (+0.1068) (+0.1080) (+0.0054) (-0.0546) (+0.0465) (+0.0165)
(2) with $\mathcal{C}_{\text{normal}}$, $\mathcal{C}_{\text{anomaly}}$ 0.9231 0.6492 0.9577 0.9106 0.9793 0.9241 0.9591 0.8072 0.9522 0.8697
(-0.0042) (-0.1325) (-0.0004) (+0.0134) (+0.0167) (+0.0672) (+0.0083) (+0.0342) (-0.0013) (+0.0783)
Results, Insights, and Future Directions. We use GPT-4o with temperature varying from 0.7 to 1.0 in multi-round synthetic generation. Table A6 presents the complete results.
LLM-generated synthetic data effectively improves AD performance. Our results show that LLM-generated synthetic data significantly enhances AD performance for several detectors. As illustrated in Figure 2(a), models like AE, ECOD, LUNAR, and VAE achieve substantial AUROC and AUPRC improvements when synthetic samples are included alongside limited real data. Notably, these models often close the gap between limited-data performance and full-data performance, demonstrating that synthetic generation can effectively compensate for data scarcity.
Performance impact varies across models. The effectiveness of synthetic generation is not consistent across all models. Methods relying on fixed geometric assumptions—such as DeepSVDD, iForest, and LOF—often degrade after augmentation (Figure 2(b)). The variance introduced by synthetic data may expand DeepSVDD’s hypersphere, perturb iForest’s isolation statistics, or blur LOF’s local-density estimates, weakening the separation between normal and anomalous points. Similarly, SO_GAAL’s adversarial training objective may become unstable as the variance may widen the definition of normal data, complicating discriminator convergence. In contrast, models like AE, ECOD, LUNAR, and VAE substantially benefit from synthetic data. Their reconstruction (AE, VAE), empirical-distribution (ECOD), or graph-aggregation (LUNAR) objectives may leverage the enriched embedding manifold, leading to more robust representations and improved detection performance. In short, synthetic generation effectively enhances detectors that learn flexible representations but can impair those reliant on fixed geometric criteria or unstable adversarial objectives.
@@ 4.3 Generating Category Descriptions for LLM-based Detectors
Results, Insights, and Future Directions. We utilize Llama 3.1, GPT-4o, and DeepSeek-V3 to generate category descriptions. We set the temperature $=0.5$ to balance the diversity and precision.
Augmented descriptions improve LLM-based AD. As shown in Table 2, incorporating category descriptions increases performance in most datasets. This suggests that the added semantic information helps LLM-based detectors discriminate anomalous samples more effectively. For example, in the “IMDB Reviews” dataset, providing richer textual representations of classes translates to noticeable gains in both metrics across LLMs.
Future Direction 4: Select Representative Samples. An effective way to refine enhanced information is to ground it in representative samples from the dataset, leveraging in-context learning of LLMs. Sampling strategies based on clustering Axiotis et al. (2024) or diversity maximization Moumoulidou et al. (2020) can identify prototype examples that guide LLMs toward producing more tailored and context-aware descriptions. By referencing these representative samples, future methods may generate more refined information that better distinguishes between normal data and anomalies, ultimately improving AD performance.
Figure 3: Model selection results across five datasets. We display the average AUROC and AUPRC of models recommended by querying each reasoning LLM five times (duplicates allowed). “Best Performance” marks the highest performance achieved by any baseline model for each dataset, while “Average Performance” denotes the mean performance across all baseline models.
@@ 5.1 Motivation
Inspired by recent research (Qin et al., 2024; Chen et al., 2024; Wei et al., 2025), LLMs offer a promising zero-shot alternative by utilizing their extensive pre-trained knowledge to analyze datasets and recommend suitable models without relying on past performance metrics. They can streamline the model selection process, reducing manual overhead and domain knowledge requirements while also improving adaptability to novel data scenarios.